What is x in a line that contains the point (2,6) and (x,15), with a slope of 3?
Answer by fractalier(6550)

We apply the slope formula
m = (y2 - y1) / (x2 - x1)
and plug in what we know
3 = (15 - 6) / (x - 2)
3(x - 2) = 9
x - 2 = 3
x = 5